1. Two yrs. ago John was three times as old as Frank. Five yrs. from now John will be twice as old as Frank. What is the sum of their ages right now?
a) 42
b) 32
c) 28
d) 21

Answer by richwmiller(9135)

(j-2)=3(f-2)
(j+5)=2(f+5)
9+23=32
b)32
Reminder:
You are limited to one problem per submission.
No more than 4 submissions daily
No similar problems.